Problem ksbonder SNF 2007-05-03 09:34:42: Wafer shifts...

gsosa at snf.stanford.edu gsosa at snf.stanford.edu
Fri May 18 16:02:41 PDT 2007


Removed ceramic pin assembly from bond head. The pin        
was siezed up in the metal collar. Had Tim from Physics lab    clean up ceramic surface. Re-assembled pin assembly and   tested. Pin now extends at least 20mm as per K/S FSE. Pin     also moves in/out smoothly.  Mario will have the system            tested this weekend.




More information about the ksbonder-pcs mailing list